# DAWN: Dynamic Frame Avatar with Non-autoregressive Diffusion Framework for talking head Video Generation

### v1

@@ 4.4 Extrapolation validation
To evaluate the extrapolation ability of our method, we conducted experiments to assess the impact of inference length on model performance using the HDTF dataset. We set the inference length to range from 40 to 600 frames in a single generation process. The results are illustrated in Table 3, suggesting that performance remains relatively stable as the inference length increases. Specifically, the image and video quality metrics, FID and FVD, are not significantly affected by the inference length. Besides, we found that the $\text{LSE}_{\text{C}}$ and $\text{LSE}_{\text{D}}$ scores tend to improve with increasing inference length while stabilizing when the length is sufficiently large. It suggests that audio with sufficient length helps the model produce more precise lip movement.
@@ 4.5 Ablation study
Ablation study on TCL strategy. Since we only use the TCL strategy on A2V-FDM, we excluded the PBNet in the evaluation process by using ground truth pose/blink to drive the video. The results are illustrated in Table 4. We separately trained A2V-FDM using either stage 1 or stage 2 of TCL strategy. According to the results, using only stage 1 causes an overall performance decrease across all metrics except FID. This is because, in the first stage, the model is trained on shorter clips. It tends to perform minor warping on the source image, which results in a lower FID score. Although minor warping can improve FID scores, it significantly diminishes the vividness of longer videos, leading to poorer FVD scores. Furthermore, training with a fixed length introduces biases into the non-autoregressive generation process, impacting the video’s smoothness and the accuracy of lip motions. This, in turn, results in worse FVD and penalties in $\text{LSE}_{\text{C}}$ and $\text{LSE}_{\text{D}}$ scores. Using only stage 2 also negatively affects the model’s performance. Compared to using only stage 1, the model achieves better $\text{LSE}_{\text{C}}$ and $\text{LSE}_{\text{D}}$ scores. This improvement is because training with longer sequences is beneficial for enhancing the precision of lip motions (Tian et al., 2024). However, the worse FID and FVD scores result from the model struggling to learn both pose/blink control and lip movement simultaneously. Using either stage 1 or stage 2 alone clearly falls behind our proposed TCL strategy, demonstrating its significant contribution to performance.
Ablation study on PBNet. We evaluate the effectiveness of the PBNet in Table 4. The term “w/o PBNet" indicates that the PBNet module was removed from the architecture, requiring the A2V-FDM to simultaneously generate pose, blink, and lip motions from the audio by itself. The results suggest an overall enhancement of all evaluation metrics with the inclusion of PBNet. This is because modeling the long-term dependency of pose and blink movements through PBNet simplifies training for the A2V-FDM. We also visualized the effectiveness of PBNet in Appendix A.2.4.
4.6 Pose/blink controllable generation
@@ A.1 Social Impact consideration
A.2 Additional experiments
A.2.1 Experiment on higher resolution and different portrait styles
We further investigate the generalization ability of our method on higher-resolution images and different portrait styles. Training DAWN on the HDTF dataset with a resolution of $256\times 256$, we tested it on multiple out-of-dataset source images featuring diverse styles, as showcased in Figure 4. The results indicate that our method yields promising outcomes in high-resolution generation and demonstrates considerable generalization ability across various image styles, including photos, paintings, anime, and sketches.
A.2.2 Comparison experiment on generation time cost
We experimented to test the time cost of video generation. Using the aforementioned methods, we generated 8-second talking head videos with the same source image and audio, then recorded the time consumption for each method. To ensure a fair comparison, we excluded the audio encoding step for all methods. The testing was performed on a single V100 16G GPU. As shown in Figure 5, our method achieves the fastest or near-fastest generation speed and requires significantly less time compared to the previous diffusion-based method, Diffused Heads.
Figure 5: The comparison experiment on generation time cost. The “*" refers to diffusion-based methods.
A.2.3 Ablation study on the local attention mechanism
@@
Table 5: Ablation study on the local attention mechanism. The “window" means the window size in the local attention operation. The “None" means we use the original attention mechanism instead.
@@
A.2.4 Ablation study on the PBNet
@@
Figure 6: The visualization of the ablation study on PBNet demonstrates different methodologies. The term “w/o PBNet" indicates “without PBNet", whereby the A2V-FDM is utilized to infer pose and blink movements. Conversely, “w PBNet" signifies the “with PBNet", which directly generates explicit pose and blink signals to control the generation of A2V-FDM.
A.3 Limitation and future works
Our work still has certain limitations. For instance, the model cannot fully comprehend the physical common sense during the generation, particularly when individuals in the portraits wearing items such as hats, helmets or headpieces. Sometimes, these items will not move along with the head, thus causing artifacts in the results. We leave the injection of these physical dependency into the model without the loss of vividness as our future work.

### v3

@@ 4.4 Extrapolation validation
To evaluate the extrapolation ability of our method, we used the HDTF dataset to assess the impact of inference length on model performance, ranging from 40 to 600 frames. FID and FVD metrics remain stable with inference length, while longer audio improves lip movement precision. This suggests that inference with sufficient length helps the model produce more precise lip movement. Furthermore, to better compare the extrapolation of other AR-based DM methods, Hallo was selected for comparison. As shown in Table 5 from the Appendix A.3, our method’s DR metric stabilizes within a certain range, in contrast to Hallo’s error accumulation at increased inference lengths.
@@ 4.5 Ablation study
Ablation study on TCL strategy. Since TCL strategy is used only on A2V-FDM, PBNet is excluded by using ground truth pose/blink for evaluation. In this section, we extend the number of training epochs based on data throughput when using single-stage training. A2V-FDM trained with either stage 1 or 2 separately shows that stage 1 decreases overall performance except for FID, due to shorter clip training causing minor warping and lower FID scores but poorer FVD. Stage 2 improves $\text{LSE}_{\text{C}}$ and $\text{LSE}_{\text{D}}$ but results in worse FID and FVD as the model struggles with simultaneous pose/blink control and lip movement. Both stages alone underperform compared to the proposed TCL strategy, underscoring its effectiveness.
Ablation study on PBNet. We evaluate the effectiveness of the PBNet in Table 4. The term "w/o PBNet" indicates that the PBNet module was removed from the architecture, requiring the A2V-FDM to simultaneously generate pose, blink, and lip motions from the audio by itself. The results suggest an overall enhancement of all evaluation metrics with the inclusion of PBNet. This is because modeling the long-term dependency of pose and blink movements through PBNet simplifies training for the A2V-FDM. We also visualized the effectiveness of PBNet in Appendix A.4.5.
@@
6 Limitation and future works
Our work still has certain limitations. For instance, the model cannot fully comprehend physical common sense during generation, particularly when individuals in portraits are wearing items such as hats, helmets, or headpieces. Sometimes, these items do not move with the head, causing artifacts in the results. We aim to inject these physical dependencies into the model without sacrificing vividness in future work. Additionally, our architecture currently requires each sub-module to be trained separately. In future work, we plan to enable joint training to reduce error propagation across modules.
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant 62171427.
@@ A.1 Social Impact consideration
A.2 Training Details
Since the HDTF dataset contains fewer videos of longer length compared to the CREMA dataset, each video in the HDTF dataset is sampled 25 times per epoch to achieve better I/O performance. The LFG is fine-tuned independently on the training datasets of CREMA and HDTF, using a checkpoint pre-trained on the VoxCeleb (Nagrani et al., 2017) dataset. We fine-tune the LFG with 400 epochs. During the training of A2V-FDM, we trained with 500 epochs for each stage. In both stages, we select the $w_{\text{lip}}=1$ for the loss function in Equation 7. We freeze the parameters of LFG when training the A2V-FDM. Although end-to-end training can prevent errors in the LFG from affecting the A2V-FDM, the main reason for not adopting end-to-end training is the difficulty in achieving stable convergence. In an end-to-end training setup, the A2V-FDM is likely to receive incorrect supervision signals, which can hinder the training process of the model and make it highly unstable. For PBNet, the MSE-based reconstruction loss can be expressed as:
$\mathcal{L}_{\text{MSE}}=\frac{1}{T}\sum_{n=1}^{N}\left(\Delta\hat{\bm{\rho}}_{n}-\Delta\bm{\rho}_{n}\right)^{2}$ (11)
where $\Delta\hat{\bm{\rho}}$ is the predicted pose/blink sequences, and $\Delta{\bm{\rho}}$ is the ground truth sequences. We also implement a KL divergence loss to constrain the latent code ${\bm{h}}$ closely approximates a standard Gaussian distribution. The adversarial loss $L_{\text{GAN}}$ is defined as:
$L_{GAN}=\arg\min_{G}\max_{D}\left(G,D\right)$ (12)
where the $G$ is our proposed PBNet and the $D$ is a discriminator implemented based on PatchGAN (Isola et al., 2016b). We use a 1D convolution based network to output the probability of each patch in the pose/blink sequence originating from real actions. The discriminator is guided by the BCE loss. In the training, the loss function of PBNet is set as:
$L=\lambda_{\text{rec}}L_{\text{rec}}+\lambda_{\text{KL}}L_{\text{KL}}+\lambda_{\text{GAN}}L_{\text{GAN}}$ (13)
where we set $\lambda_{\text{rec}}=1$, $\lambda_{\text{GAN}}=0.6$. During the training process, we observed that $L_{\text{KL}}$ rapidly converges to approximately zero at the beginning. This convergence impairs the latent state’s ability to retain effective information, causing the model to depend almost entirely on the decoder’s predictive capability for completing the fixed fitting task from audio to pose/blink. To ensure diversity in generation, we implemented a method of progressively increasing the $\lambda_{\text{KL}}$ in training PBNet. The PBNet was trained over 1600 epochs. In the initial 400 epochs, we did not apply the $L_{\text{KL}}$ constraint to the latent code., which helped the model develop basic sequence reconstruction capabilities in the early training phase. From that point until the end of training, $\lambda_{\text{KL}}$ was gradually increased to 0.01.
@@ A.2 Training Details
A.3 Additional experiments
Table 5: The comparison experiment of error accumulation with Hallo. "Inference length" refers to the number of frames generated in a single inference process.
Inference length DAWN Hallo
$\text{DR}_{25}$ $\text{DR}_{50}$ $\text{DR}_{25}$ $\text{DR}_{50}$
200 0.253 0.043 0.214 0.094
400 0.208 0.164 0.279 0.161
600 0.152 0.152 0.422 0.332
A.3.1 Extrapolation Comparison
To highlight the advantages of NAR generation, we selected an image-based diffusion method, Hallo, for comparison, due to its representative structure similar to most of the latest AR or SAR methods. We evaluated the degradation rate for both our model and Hallo across inference lengths of 200, 400, and 600, as detailed in Table 5. The results indicate that, in our method, both the $\text{DR}_{25}$ and $\text{DR}_{50}$ metrics remain stable as inference length increases. In contrast, Hallo’s $\text{DR}_{25}$ and $\text{DR}_{50}$ metrics significantly increase with longer inference lengths, suggesting notable error accumulation in the image space. This demonstrates that image-based AR methods, such as Hallo, struggle with error accumulation in longer tasks, while our model exhibits superior extrapolation capabilities and more consistent performance with extended video sequences. These findings underscore the superiority of NAR methods compared to AR and SAR alternatives.
A.4 User study
The user study evaluates generated videos across four dimensions: 1) L-Sync: Lip-audio synchronization; 2) O-Nat: The overall naturalness of the generated talking head ; 3) M-Viv: The vividness of the head movements; 4) V-Qual: The overall video quality (e.g., presence of artifacts or abnormal color blocks). We generated 10 test videos per method, with 23 participants scoring each on a 1-5 scale. Users disregarded resolution and cropping when rating. As shown in Table 6, our method surpasses existing approaches in lip synchronization, naturalness, and head movement vividness, and is comparable to EchoMimic in video quality.
Table 6: User study.
Method L-Sync O-Nat M-Viv V-Qual
Audio2Head 3.87 3.67 3.28 3.66
Sadtalker 4.23 3.13 2.81 4.14
DreamTalk 4.38 3.89 3.41 4.42
Hallo 4.40 3.76 3.89 4.03
EchoMimic 4.45 4.30 4.06 4.53
DAWN(ours) 4.57 4.41 4.43 4.51
A.4.1 Pose/blink controllable generation
A.4.2 Experiment on higher resolution and different portrait styles
We further investigate the generalization ability of our method on higher-resolution images and different portrait styles. We trained DAWN on the HDTF dataset with a resolution of $256\times 256$. Then we evaluated our higher-resolution model on the HDTF test set. The quantitative results are illustrated in Table 7. The results show that our method still maintains strong competitiveness compared to the latest talking head models. We also test our method on multiple out-of-dataset source images featuring diverse styles, as showcased in Figure 3. The results indicate that our method yields promising outcomes in high-resolution generation and demonstrates considerable generalization ability across various image styles, including photos, paintings, anime, and sketches.
A.4.3 Comparison experiment on generation time cost
We experimented to test the time cost of video generation. We generated 8-second talking head videos with the same source image and audio, then recorded the time consumption for each method. To ensure a fair comparison, we excluded the audio encoding step for all methods. The testing was performed on a single V100 16G GPU. As shown in Figure 5, our method achieves the fastest or near-fastest generation speed and requires significantly less time compared to the previous diffusion-based methods, Diffused Heads, Hallo, and EchoMimic.
Table 7: Quantitative study on different resolutions. The "128" indicates our method is trained on a $128\times 128$ resolution, and the "256" indicates training on a $256\times 256$ resolution.
[…]
GT - - - 7.95 7.33 1 0.267 0.75
DAWN (128) 9.60 60.34 95.64 6.71 7.94 0.790 0.281 0.86
DAWN (256) 11.80 68.07 105.20 7.20 7.80 0.791 0.278 0.73
Figure 5: The comparison experiment on generation time cost. The Diffused Heads, Hallo, EchoMimic are existing diffusion-based methods.
A.4.4 Ablation study on the local attention mechanism
@@
Table 8: Ablation study on the local attention mechanism. The "window" means the window size in the local attention operation. The "None" means we use the original attention mechanism instead.
@@
A.4.5 Ablation study on the PBNet
@@
Figure 6: The visualization of the ablation study on PBNet demonstrates different methodologies. The term "w/o PBNet" indicates "without PBNet", whereby the A2V-FDM is utilized to infer pose and blink movements. Conversely, "w PBNet" signifies the "with PBNet", which directly generates explicit pose and blink signals to control the generation of A2V-FDM.